OpenStudy (anonymous):

Find the volume of the solid enclosed by the planes 4x+2y+z=10, y=3x, z=0, x=0

OpenStudy (anonymous):

I am just having trouble sketching the region.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

This is a double integral question btw.

OpenStudy (anonymous):

This isn't algebra!!! :P .

OpenStudy (luigi0210):

@Dido525 I can provide you with a graph? I'm not sure if it's good though >.<

OpenStudy (anonymous):

I just want to know how you would sketch these kinds of graphs >.< .

OpenStudy (anonymous):

I can solve the double intergal fine :/ .

OpenStudy (anonymous):

So yeah sketch it please :) .

OpenStudy (luigi0210):

Well it's not really a sketch :3

OpenStudy (anonymous):

That just makes it look worse >.< .

OpenStudy (anonymous):

Go to the last page. How did they sketch that?

OpenStudy (anonymous):

Last example*

OpenStudy (anonymous):

Example 4.

OpenStudy (luigi0210):

Btw hi @Dido525 I'm Luigi, I'm not even a Calculus one student :3

OpenStudy (anonymous):

Ohh okay XD . Thanks anyways :) .

OpenStudy (psymon):

Lol, Luigi summons me here when I cant do calc 3 yet. Dude, calc 2 and under @Luigi0210 !!

OpenStudy (luigi0210):

Sorry had no idea >.,

OpenStudy (anonymous):

Ahh Calc II. Good times :) .

OpenStudy (psymon):

Im doing self-study on calc 3 xD Im barely at intro to quadric surfaces xD

OpenStudy (anonymous):

We did that in calc II :P .

OpenStudy (psymon):

Different places different things in each calc class xD I did sequences and series + intro to DEs in calc 2.

OpenStudy (anonymous):

Same :P .

OpenStudy (psymon):

Yeah. Other places have a calc 4 xD I wonder if that means we move faster than them here >.<

OpenStudy (dumbcow):

for 3-d graphing http://graphcalc.com/index.shtml

OpenStudy (dan815):

hi

OpenStudy (anonymous):

Hi.

OpenStudy (dan815):

do you still need help?

OpenStudy (anonymous):

Yes.

OpenStudy (dan815):

ok lemme read

OpenStudy (dan815):

Find the volume of the solid enclosed by the planes 4x+2y+z=10, y=3x, z=0, x=0

OpenStudy (dan815):

so u got plane with normal 4,2,1 and that stuff

OpenStudy (dan815):

u can write z in tersm of x and y

OpenStudy (anonymous):

I get HOW to find the volume. But I don't know how they sketched the plane.

OpenStudy (dan815):

do a triple integral in terms of z=f(x,y) then y=f(x) and x from 0 to w/e

OpenStudy (dan815):

oh okay

OpenStudy (anonymous):

No no. I am not at triple intergals yet.

OpenStudy (dan815):

sketching is easy pzy

OpenStudy (dan815):

lets start with the plane

OpenStudy (anonymous):

But I can do this with double intergals. I just need to see the plot of this.

OpenStudy (dan815):



OpenStudy (anonymous):

Yep.

OpenStudy (dan815):

ya sure u can do it as a double integral.. but ur still gonna use a triple integral simplified to a double integral format

OpenStudy (dan815):

what did they teach you like top function - bot function and then double integral over the domain?

OpenStudy (anonymous):

That's fine. I just need a sketch.

OpenStudy (anonymous):

yeah Kinda.

OpenStudy (dan815):